If a striaught line intercepts with axes of equal magnitude but of opposite signs and passes through point P(3,4). Find the equation of the straight line and hence determine its slope.

Grade:6

1 Answers

Arun
6 years ago
let assume the line is L :
x/ a + y/ –a  = 1
pass this line from (3, 4)
(3 – 4) / a = 1
a = –1
 
hence equation of line L:
x/ –1  + y/ 1  = 1
 – x + y = 1
 
and slope  will be m =1